a farmer has 1 hectare of land. one half of the land was used for planting oranges, 1/3 of the remainder was used for planting mangos while plantain was planted on the rest. Express the area of land used for mangoes as fraction of that used for plantain
Answer by Boreal(15235)

1/2 used for oranges
1/3 (1/2)=1/6 used for mangos.
The rest is 1/3
mangos/plantain=(1/6)/(1/3)=1/2.
1/3 of the remainder was for mangoes and 2/3 for plantain, so 1/2.